# LLM Jailbreak Attack versus Defense Techniques - A Comprehensive Study

### v1

@@ 7 Conclusions
In this work, we present the first comprehensive assessment of existing attack and defense strategies in the context of LLM security. Additionally, we contribute to the field by releasing the first framework specifically designed for assessing the robustness of LLMs against various threats. We selected nine attacks and seven defensive mechanisms from existing literature and software libraries for our analysis. Our experimentation, conducted on three distinct models, reveals that Template methods are notably effective, with an average of 78 templates identified as critical for thwarting jailbreak attempts. Regarding Generative methods, GPTFuzz emerged as the most efficacious. Our investigation into question categorization demonstrated that all three models exhibit enhanced resilience against queries related to harmful content and illegal activities. However, our analysis of current defensive measures indicates a general ineffectiveness, with Bergeron showing comparatively better performance. We highlight the necessity of establishing a uniform baseline for jailbreak detection, as existing defenses employ varied methodologies, and the need to develop better defense techniques. Additionally, our study observed the impact of using the ’[/INST]’ marker in the Llama model. Looking forward, we aim to continuously incorporate evolving attacks and defenses into our framework, thereby providing a dynamic overview of the field’s progression.
@@ 9 Ethical Considerations and Disclaimer
Throughout the course of our investigation, we have steadfastly avoided engaging in any activities that involve the creation, dissemination, or promotion of content that could be deemed malicious or harmful. Our approach to research is guided by a strong ethical compass, ensuring that our methodologies and findings do not contribute to the proliferation of such content.
@@ Appendix A APPENDIX
Table 7: The attack results of Vicuna, the top three best attacks in terms of ASR and efficiency are highlighted
@@
Table 8: The attack results of Llama, the top three best attacks in terms of ASR and efficiency are highlighted. Although the ASR of the Parameter is slightly lower than that of the Pair, its significantly higher efficiency positions the Parameter as the better choice.
@@
Aegis 0.00% (0/805) 0.00% (0/56) 0.00% (0/0) 0.00% (0/75) 0.00% (0/79) 0.00% (0/119) 0.00% (0/975) 0.00% (0/767) 0.00% (0/1559) 0.00% (0/0) 0.00%
llm-guard 99.13% (798/805) 0.00% (0/56) 0.00% (0/0) 0.00% (0/75) 100% (79/79) 100.00% (119/119) 33.33% (325/975) 9.51% (73/767) 97.62% (1522/1559) 0.00% (0/0) 37.82%
smooth-llm 93.79% (755/805) 53.57% (30/56) 0.00% (0/0) 37.33% (28/75) 78.48% (62/79) 77.31% (92/119) 12.82% (125/975) 35.20% (270/767) 10.07% (157/1559) 0.00% (0/0) 27.91%
Baseline-defense 69.07% (556/805) 0.00% (0/56) 0.00% (0/0) 10.66% (8/75) 89.87% (71/79) 94.11% (112/119) 33.33% (325/975) 3.12% (24/767) 82.16% (1281/1559) 0.00% (0/0) 34.80%
RA-LLM 88.45% (712/805) 76.78% (43/56) 0.00% (0/0) 60.00% (45/75) 67.08% (53/79) 59.66% (71/119) 15.89% (155/975) 57.88% (444/767) 5.83% (91/1559) 0.00% (0/0) 29.59%
Bergeron 98.51% (793/805) 12.5% (7/56) 0.00% (0/0) 5.33% (4/75) 25.31% (20/79) 22.68% (27/119) 5.74% (56/975) 7.95% (61/767) 7.24% (113/1559) 0.00% (0/0) 8.25%
ModerationAPI 99.63% (802/805) 100% (56/56) 0.00% (0/0) 77.33% (58/75) 98.73% (78/79) 99.15% (118/119) 88.00% (858/975) 88.78% (681/767) 96.72% (1508/1559) 0.00% (0/0) 60.96%
@@
Model Smooth Rallm Bergeron

### v2

@@ 7 Conclusions
In this work, we present the first comprehensive assessment of existing attack and defense strategies in the context of LLM security. Additionally, we contribute to the field by releasing the first framework specifically designed for assessing the robustness of LLMs against various threats. We selected nine attacks and seven defensive mechanisms from existing literature and software libraries for our analysis. Our experimentation, conducted on three distinct models, reveals that Template methods are notably effective, with 78 templates technique identified as the most powerful one. Regarding Generative methods, GPTFuzz emerged as the most effective given the experiment budget. Our investigation into question categorization demonstrated that all three models exhibit enhanced resilience against queries related to unlawful practice, harmful content and illegal activities. However, our analysis of current defensive measures indicates a general ineffectiveness, with Bergeron showing comparatively better performance. We highlight the necessity of establishing a uniform baseline for jailbreak detection, as existing defenses employ varied methodologies, and the need to develop better defense techniques. Additionally, our study observed the impact of using the ’[/INST]’ marker in the Llama model. Looking forward, we aim to continuously incorporate evolving attacks and defenses into our framework, thereby providing a dynamic overview of the field’s progression.
@@ 9 Ethical Considerations and Disclaimer
Aware of the fine line between knowledge advancement and safety assurance, we introduced measures like limiting the length of potentially malicious responses in our dataset. This method aims to support evaluation and learning without revealing practical information prone to misuse. We emphasize our dedication to ethical practices by actively reducing the risk of spreading harmful content.
@@ Appendix A APPENDIX
A.1 Analysis of Categorization of Attack Techniques
Figure 5: This graph assesses the pros and cons of three attack categories across five dimensions.
Our empirical analysis and experimental results identified five metrics for assessing the advantages and disadvantages of various attack techniques, as shown in Figure 5.
The criterion of Complexity measures the intrinsic algorithmic challenge posed by each method. Notably, the Generative approach is identified as the most complex, attributed to its sophisticated algorithmic underpinnings. This is followed by the Training Gaps method, which demands substantial insight into the model’s operation for effective application.
The dimension of Specificity evaluates whether an attack is tailor-made for a particular model. Given that Training Gaps are dependent upon the unique safety training protocols of each model, they inherently exhibit the highest specificity. Subsequently, the Template-Based method, often crafted for specific model types (e.g., the GPT series), ranks next in specificity.
In terms of Ease of Use, the Template-Based approach emerges as the most user-friendly, attributed to its pre-designed nature, thereby facilitating immediate application. The Training Gaps method follows, offering relatively straightforward deployment when contrasted with the more complex Generative approach.
Regarding Ease of Fix, Template-Based attacks, due to their predefined structure, allow for direct incorporation into safety training protocols, simplifying mitigation efforts. Similarly, addressing vulnerabilities exposed by Training Gaps is comparatively easier.
Lastly, the criterion of Running Cost reveals that Generative techniques, due to their intensive iteration and deployment requirements, incur the highest expenses. The Template-Based method, necessitating the processing of extensive prompts, ranks second, surpassing Training Gaps in terms of token processing demands.
A.2 Analysis of Categorization of Defense Techniques
Figure 6: This graph shows a comparative analysis of defense categories across four dimensions.
Similarly, we designed four key metrics for evaluating the pros and cons of different defense techniques.
Autonomy assesses the degree to which a model depends on external resources for detection. Self-Processing exhibits the highest autonomy, followed by Input Permutation—this method necessitates further processing of the input and multiple generations with the LLM. Conversely, the Additional Helper method employs an external algorithm or model for decision-making.
Running Cost evaluates the operational expenses; Input Permutation is notably resource-intensive due to significant input modifications and subsequent model verifications, making it more expensive than Additional Helper, followed by Self-Processing. However, Rain (Li et al., 2023c) of Self-Processing category presents an outlier, incurring prolonged processing times due to its auto-regressive input examination.
Adaptability assesses the capability of the approach to adjust to emerging attack vectors. In light of the continuous progress and enhancement of detection algorithms in LLMs, the Additional Helper strategy, typically involving the integration of a further sophisticated model or algorithm, gains from persistent updates. Likewise, the Self-Processing approach benefits from this adaptability.
Lastly, Comprehensiveness measures a defense mechanism’s capability to generalize across attack types. Employing Additional Helper specializing in identifying malicious inputs ranks highest in effectiveness, followed by Self-Processing, which is inherently limited by the model’s capabilities, and input permutation, which usually disrupts the attacks utilize the embedded gradient information such as GCG (Zou et al., 2023).
A.3 Customized Evaluator
With a dataset encompassing 199,427 responses, the task of identifying malicious content posed significant challenges due to the impracticality of manual review and the absence of standardized evaluation methods for malicious responses—a gap in current research.
Existing strategies for addressing this issue vary. Zou et al. (2023) utilized a set of common refusal patterns, such as "I am sorry" and "I cannot", to automate the identification of non-compliant responses. Yu et al. (2023) and Huang et al. (2024) focused on enhancing machine learning models, specifically RoBERTa and BERT-BASE-CASED models, respectively. Additionally, approaches leveraging GPT-4 for attack analysis were explored by Chao et al. (2023) and Mehrotra et al. (2023).
To ensure a rigorous analysis of our dataset, we selected a random sample of 1,068 responses, aiming for a 95% confidence level with a 3% margin of error. This sample underwent manual annotation to assess each response’s relevance to the inquiry and its alignment with providing a malicious answer. An example of this annotation process is documented in Figure 9, and a comparative analysis of the effectiveness of different models is provided in Table 5.
Reflecting on the reported efficacy of the fine-tuned RoBERTa model by Yu et al. (2023), we chose to refine this model further utilizing our manually annotated dataset, accessible on HuggingFace (fine tuned, 2024). The fine-tuning protocol involved a batch size of 5, three training epochs, a learning rate of $2\times 10^{-5}$, application of the Adam optimizer, and a linear rate decay complemented by a warm-up phase covering 10% of the training duration. Post-labeling, an additional round of random sampling was conducted for manual verification to ascertain the accuracy and reliability of our findings.
Figure 7: The graph depicts the appearance of jailbreak prompts corresponding to various attack techniques.
@@ A.3 Customized Evaluator
Table 7: The attack results of Vicuna, the top three best attacks in terms of ASR and efficiency are highlighted.
@@
Table 8: The attack results of Llama, the top three best attacks in terms of ASR and efficiency are highlighted. Although the ASR of the Parameter is slightly lower than that of the Pair, its significantly higher efficiency positions the Parameter as the better choice. The GCG on LLama is configured to perform 500 iterations. This setting is based on empirical evidence indicating that 75 iterations fail to produce jailbreak outcomes for the majority of queries processed by GCG on Llama, as loss shown in Figure 4. Averagely, 204 iterations are required to jailbreak a question in our dataset on Llama. Despite this increasing, the universal methods except for DeepInception still demonstrate better performance.
@@
Aegis 0.00% (0/805) 0.00% (0/56) 0.00% (0/0) 0.00% (0/75) 0.00% (0/79) 0.00% (0/119) 0.00% (0/975) 0.00% (0/767) 0.00% (0/1559) 0.00% (0/54) 0.00%
llm-guard 99.13% (798/805) 0.00% (0/56) 0.00% (0/0) 0.00% (0/75) 100.00% (79/79) 100.00% (119/119) 33.33% (325/975) 9.51% (73/767) 97.62% (1522/1559) 96.29% (52/54) 48.52%
smooth-llm 93.79% (755/805) 53.57% (30/56) 0.00% (0/0) 37.33% (28/75) 78.48% (62/79) 77.31% (92/119) 12.82% (125/975) 35.20% (270/767) 10.07% (157/1559) 0.00% (0/54) 33.86%
Baseline-defense 69.07% (556/805) 0.00% (0/56) 0.00% (0/0) 10.66% (8/75) 89.87% (71/79) 94.11% (112/119) 33.33% (325/975) 3.12% (24/767) 82.16% (1281/1559) 0.00% (0/54) 34.80%
RA-LLM 88.45% (712/805) 76.78% (43/56) 0.00% (0/0) 60.00% (45/75) 67.08% (53/79) 59.66% (71/119) 15.89% (155/975) 57.88% (444/767) 5.83% (91/1559) 0.00% (0/54) 38.12%
Bergeron 98.51% (793/805) 12.5% (7/56) 0.00% (0/0) 5.33% (4/75) 25.31% (20/79) 22.68% (27/119) 5.74% (56/975) 7.95% (61/767) 7.24% (113/1559) 10.52% (6/54) 10.80%
ModerationAPI 99.63% (802/805) 100% (56/56) 0.00% (0/0) 77.33% (58/75) 98.73% (78/79) 99.15% (118/119) 88.00% (858/975) 88.78% (681/767) 96.72% (1508/1559) 87.03% (47/54) 81.74%
@@
Model Smooth GRQ Rallm GRQ Bergeron GRQ